# thanks: Contributed equallythanks: Contributed equally

### v3

@@ III.1 CLFs via long-ranged repulsive torques
We first consider the case of $\chi_{t}=0$, thus repulsions are only short-ranged that avoid overlap of the particles in the system via $\mathbf{F}_{i}$. The phase diagram delimiting polar flocks (or CLF), bands, and disordered states are shown in Figure 1(a). The states are delimited via the mean global polarization of the system at the steady state, $\langle|\mathbf{P}|\rangle_{ss}$, where the polarization $\mathbf{P}$ is defined as the average of sum of the orientation of all the particles:
@@
On the $(\rho,\chi_{r})$ plane (Fig. 1(a))), we see that low density flocks are sustained, that are eventually destabilized at sufficiently high repulsion and densities. These flocks have evidently a liquid-like spatial structure (Fig. 1(b); see also Fig. 3(d)) [25, 36, 37]. The density field $\rho$ is thus not homogeneous across the system (see Fig. 1(c)). We note that $\tau={b}/{v_{0}}$ is the propulsion time scale. In addition, we define $\tau_{r}=b^{4}/\chi_{r}$, which sets the time scale for deterministic rotation of the particle due to phoretic interactions.
The phenomenology of the CLF can also be understood by a comparison of the relevant length scales of the problem. First, we write down natural length scales that arise out of the equations of motion (1). They are:
$l_{r}$ $=\frac{b\,\tau}{\tau_{r}}=\frac{\chi_{r}}{b^{2}v_{0}},\qquad\quad l_{t}=\frac{\chi_{t}}{b^{3}v_{0}}.$ (6)
Here, $l_{r}$ is a deterministic length scale constructed by multiplying the radius $b$ of the particle with the ratio of propulsion time scale $\tau$ and time scale $\tau_{r}$ which controls rotations of the particle orientations from chemical interactions between particles. On the other hand, $l_{t}$ is a deterministic translational length, which measures ratio of translation due to chemical interactions between particles and one-body propulsion. From the phase diagram, we see that there are separate length scales that arise as a consequence of the collective many-body dynamics. To this end, let us construct a sliding time scale $\tau_{sl}$ and length scale $l_{sl}$. The time-scale $\tau_{sl}$ is time during which pair colloids slide a distance $l_{sl}$ whilst rotating away. Similarly, $l_{sl}$ represents the typical length the pair colloids would co-move upon collision before rotating away. The sliding time $\tau_{sl}$ and length $l_{sl}$ can be written as:
$\tau_{sl}=\frac{b^{2}}{v_{0}l_{r}},\qquad l_{sl}=v_{0}\tau_{sl}=\frac{v_{0}b^{4}}{\chi_{r}}.$ (7)
We next use these to understand the mechanism of formation of CLF.
Let us denote $l_{-}$ as the smallest value of $l_{r}$ to sustain the CLF (corresponding to the lowest transition line in Fig.1a), and $l_{+}$ as the largest value of $l_{r}$ above which the CLF is not sustained. $l_{+}$ is density dependent, as we describe below. The mechanism for the CLF can then be understood as a competition between $l_{r}$ and $_{+}, ll_{-}$; this is illustrated in Fig. 2 (annotated accordingly in the phase diagram of Fig.1(a)). Equivalently, the mechanism can be understood in terms of sliding time scale $\tau_{sl}$ along with $\tau_{sl}^{-}$ and $\tau_{sl}^{+}$. We now use the schematic of Fig. 2 to explain a series of transitions from disorder-to-order-to-disordered phases. When $l_{r}<l_{-}$ (corresponding to $\tau_{sl}>\tau_{sl}^{-}$), we have a sub-deterministic sensing of the chemicals; here the orientational changes of the colloids are too slow compared to the free movement time, pair colloids slide excessively long, hence local polar order cannot be sustained. See panel (I) of Fig. 2. Increasing $l_{r}$, in the regime of $l_{-}<l_{r}<l_{+}$ (equivalently $\tau_{sl}^{+}<\tau_{sl}<\tau_{sl}^{-}$), we have that the chemical response time fast enough such that local polar order is sustained, but still sufficiently slow such that the forward-backward symmetry at the pair collision level is broken. See panel (II) of Fig. 2. Finally for $l_{r}>l_{+}$ (equivalently $\tau_{sl}<\tau^{+}_{sl}$), the chemical response is sufficiently quick such that there is no symmetry breaking at the pair collision level. See panel (III) of Fig. 2. We note that from simulations at $\rho\approx 0.1$, we find these values to be $l_{sl}^{+}\sim 8.3b$ and $l_{sl}^{-}\sim 100b$. Here $l_{sl}^{+}=v_{0}\tau^{+}_{sl}$ and $l_{sl}^{-}=v_{0}\tau^{-}_{sl}$ (note that in our notation choice $l_{sl}^{-}>l_{sl}^{+}$). We conclude that a pair of colloids have to slide at least a distance $l_{sl}^{+}$, before deterministically turning away, as the underlying mechanism for the CLF.
III.2 Effect of long-ranged translational repulsion
@@
We now switch on $\chi_{t}$ and ask how the properties of the flock differ. On the $(l_{r}/b,l_{t}/b)$ plane, we find that there is a clear region of flocking for non-zero $\chi_{t}$ (see Figure 3(a)). We further find that the flocks have crystalline order (see Figure 3(c),(e)). Such crystalline flocks in repulsive systems have been recently reported [26]. Similarly to the CLF, the CCF destabilized by sufficiently high rotational torques. In addition, $l_{+}$ now is linear in $_{t}. l+l_{+}$ being larger for a positive $l_{t}$, this thus renders a shorter sliding length for pair colloids during the CCF formation. The net result is thus a shorter sliding length compared to the CLF, which we can attribute to the difference between crystalline and liquid structures. We note that if this repulsion is too strong, the CCF is destabilized to a disordered gas (not shown here). The most generic form of the transition lines from the phase diagrams can be written as
$l_{r}<l_{+},\qquad\frac{l_{+}}{b}=\frac{c_{1}}{\rho}+\frac{\beta_{1}l_{t}}{b}$ (8)
which is presented in Figures (1)(a) and (3)(a). We obtain $\beta_{1}$ and $c_{1}$ by first fitting the boundary line in Fig. 3(a). The slope gives $\beta_{1}\approx 0.167$, whilst the $x$-intercept gives $c_{1}\approx 0.013$. This $c_{1}$ then independently fixes the boundary on Fig. 1(a). We show in section IV how this can be derived from a continuum model of interacting density and polarization fields. The $\frac{c_{1}}{\rho}$ term inhibits sliding at high densities. We note that for densities above a critical density $\rho_{c}=(\frac{c_{1}}{b})^{2/3}\approx 0.05$, $l_{sl}$ always exceeds the mean free path $\frac{1}{\sqrt{\rho}}$ so that flock formation can happen uninhibited. Thus, the destabilization at higher densities occurs as $l_{sl}^{-}$ is approached due to the dominance of the instantaneous chemical sensing relative to the excluded volume repulsions (though in this regime collisions are more frequent). For $\chi_{t}>0$, the second term further suppresses the sliding length due to long-ranged translational repulsion. Sliding over shorter distances effectively localizes particles in space and the effect is a crystalline structure.
We note that the flock destabilization mechanism here is independent of the contribution of noise. Though we have explicit noise terms in the updates in Eq.(1) for the sake of completion, they are set to be negligibly small (see Table 1 in Appendix). Indeed, if we were to define length scales $l_{\eta}=\frac{D_{r}b^{2}}{v_{0}}$ and $l_{\xi}=\frac{D_{t}}{v_{0}}$, using the values used of $D_{r}\approx 10^{-{6}}$ and $D_{t}\approx 10^{-{8}}$, we find that $l_{\eta}\sim 10^{-4}b$ and $l_{\xi}\sim 10^{-6}b$, thus negligible compared to all other length scales in the problem. We have repeated all the simulations presented in Fig. (1) with $D_{r}=D_{t}=0$, and have found the results to be identical. Thus, as opposed to the conventional picture of noise induced inhibition of flock formation [4, 10, 21, 26] (i.e. deterministic alignment interactions compete with randomness in the same interaction channel), our mechanism reports a competition between a deterministic length scale due to sensing of chemical gradients $l_{r}$ (also $l_{t}$) and other deterministic length scales of the problem.
III.3 Difference between CCF and CLF
@@
The difference in spatial structure between the CLF and CCF is readily distinguished by the pair correlation function $g(r)=\frac{1}{N}\sum_{i,j}\delta(r-|\mathbf{r}_{i}-\mathbf{r}_{j}|)$. This is plotted in Fig. (3)(f), distinguishing the solid and liquid signatures [25]. A closely related quantity [38] is the polar pair correlation $g(r,\phi)$ which explicitly depends on the orientations of the colloids. Here, $\cos(\phi_{ij})=\mathbf{e}_{i}\cdot(\hat{r}_{i}-\hat{r}_{j})$[38]. For the CLF, we see that there are quadrants where neighbours are more likely to be found (3(d)), as opposed to the hexatic order of the CCF (3(e)). In the case of CCF, the intensity (proportional to the probability of finding the particle from the reference - see Appendix B) in the well-localized hexagonal positions of the crystal are substantially higher than in the CLF, which are instead less intense with a greater relative spread. The quadrant-like distribution indicate the repulsive torque-induced turn away of the particles in the forward and backward direction of the reference particle.
To further support the robustness of the mechanism, we report CLFs arising from repulsive torques with the addition of long-ranged attractive interactions, thus $\chi_{t}<0$. We find that this phenomenology is maintained for $l_{t}\in[0,-0.01]$, before transitioning to a disordered gas for larger attractive translational interactions; this is shown in Figure 3(b). Putatively, the clumping tendency in this range is balanced by the excluded volume repulsions.
@@
Figure 4: Phase diagrams in terms of (left) global hexatic order ($\psi_{6}$) and (right) density variance. The panel on flocktop shows results for the case (a), which is $\chi_{t}=0$ (CLF), whereas the bottom panel shows results for the case (b), which is $\chi_{t}>0$ (CCF). Dotted curves indicate flocking order-disorder transition line. See text for details.
III.4 Hexatic order and density variance
@@
Here, $\theta_{ij}$ is the angle between particle $i$ and particle $j$, and $N^{n}_{i}$} is the number of neighbors for particle $i$ ($N^{n}\approx 6$ for all data points). The observable $\psi_{6}$ is then averaged over $1000$ different particle configurations at the steady-state. We plot $\psi_{6}$ for the two cases ($\chi_{t}=0$ and $\chi_{t}>0$) in the left panel of Fig. 4. The transition lines from the phase diagrams in Figures 1(a) and 3(a) are also shown as dotted lines for guide to the eye. In the case of $\chi_{t}=0$ (Fig. 4(a) left), the hexatic order exists in the flocking phase, only for a very high density (area fraction $\rho\pi b^{2}>0.6$. Whereas, in case of $\chi_{t}>0$ (Fig. 4(b) left), with $\rho\pi b^{2}=0.36$, the system becomes highly hexatic in the flocking phase. However, near the transition line, where density bands are prone to occur, the hexatic order gradually decreases.
@@
Next, density variance for the two cases ($\chi_{t}=0$ and $\chi_{t}>0$) are shown in the right panel of Fig. 4. The density variance [15] is calculated from the density distribution for a specific parameter set shown in the phase diagram. For local density measurements, the Voronoi tessellation method has been implemented for a spatial configuration of particles. We calculate local density $\rho_{l}=\frac{\pi b^{2}}{A_{l}}$, where $A_{l}$ is local area assigned to each particle. Then we compute the density variance as $\langle\rho_{l}^{2}\rangle-\langle\rho_{l}\rangle^{2}$ (spatial average) and finally averaged over $1000$ configurations. In both cases, the density variance is much lower in the flocking state. However, it shows an increase near the transition region. The density variance is high for the high area fraction ($\rho\pi b^{2}$) and in the disordered phase, for the case $\chi_{t}=0$ (Fig. 4(a) right panel). For $\chi_{t}>0$ case, with low $\chi_{t}$ values, the variance is more prominent, shown in Fig. 4(b) (right panel).
III.5 CCFs as solutions to equation of motion
Our choice of EOM in (1) is also unique in that it can be shown to support crystalline flocks as its dynamical fixed points. To see this, we firstly note that from the EOM the fixed points of $\theta_{i}$ are obtained by setting $\Omega_{i}=0$, giving
$\theta_{i}^{*}=\tan^{-1}\Big{(}\frac{\sum_{j\neq i}(y_{i}-y_{j})\prod_{k\neq i,j}|r_{i}-r_{k}|^{3}}{\sum_{j\neq i}(x_{i}-x_{j})\prod_{k\neq i,j}|r_{i}-r_{k}|^{3}}\Big{)}$ (10)
The hexagonal lattice topology of CCF makes (10) tractable. Firstly, let us consider a hexagonal lattice topology of distance $2b^{\prime}$ aligned along the direction of the flock $\theta_{i}$ for the $i$th reference particle. To compute the sums in (10) one needs to evaluate the sums
$\sum_{j}r_{i}-r_{j}=-b^{\prime}\sum_{k=0}^{K}\sum_{n=1}^{6}(1+k)\begin{pmatrix}\cos\Big{(}(n-1)\frac{\pi}{3}-\theta_{i}\Big{)}\\ \sin\Big{(}(n-1)\frac{\pi}{3}-\theta_{i}\Big{)}\ \end{pmatrix}$ (11)
where $K$ is the upper bound on the radial sums performed on the lattice; in our case $K=\frac{50}{b^{\prime}}-1$. For $\prod_{k\neq i,j}|r_{i}-r_{k}|^{3}$, we can evaluate this to be $\prod_{k}^{K}((1+k)b^{\prime})^{z-1}\delta_{k,j\in\{k\}}((1+k)b^{\prime})^{z}(1-\delta_{k,j\in\{k\}})$, thus raised to power coordination number $z$ minus one for the radial coordinate containing the $j$th particle (denoted by $j\in\{k\}$), and to the coordination number for all other radial coordinates. We recognize immediately that the symmetry of the lattice structure means that this depends on only the location of the “left out” particle $j$ and the reference particle $i$. Without loss of generality, we label this product $P_{ij}$. The sums over $n$ for each component in (11) can be readily computed by applying the double angle formulae and symmetries of the sine and cosine. The overall result is thus
$\sum_{j}r_{i}-r_{j}=-b^{\prime}\begin{pmatrix}\cos(\theta_{i})\\ \sin(\theta_{i})\end{pmatrix}S_{K}$ (12)
where we have defined $S_{K}=\sum_{k=0}^{K}(1+k)$ (not needed to be evaluated explicitly). Thus the fixed point angle of (10) simply reads
$\theta_{i}^{*}=\tan^{-1}\Big{(}\frac{\sum_{j}(y_{i}-y_{j})P_{ij}}{\sum_{j}(x_{i}-x_{j})P_{ij}}\Big{)}$ (13)
which we see to be the self-consistent $\theta_{i}^{*}=\theta_{i}$ as required. Further, taking ratio of the coordinates of the positional dynamics at the fixed point, we see that $\frac{\dot{y}_{i}}{\dot{x}_{i}}=\tan(\theta_{i})$ as required. We thus conclude that our particular choice of chemical interactions of (4) supports a crystalline flock. For fluid flocks the result is less trivial since various sums and products need to be weighted by an appropriate distribution function. We nevertheless see that enhanced fluctuations of $\theta_{i}^{*}$ is to be expected, since any approximate evaluation of (10) will contain various moments of each components of (11).
III.6 Order-disorder phase transition
Figure 5: Magnitude of the polar order parameter $|\mathbf{P}|$ versus $l_{r}/b$ in panel (a) for the case of CLF. Here $l_{t}=0$. In panel (b) we show, for the case of CCF, the variation of the order parameters as a function of $l_{t}/b$. Here, we have fixed $l_{r}=0.55$. Different symbols indicates different area fraction $\phi=\rho\pi b^{2}$.

### v4

@@ III CLF via chemo-repulsive torques
We first consider the case of with $\Lambda_{t}=0$ (or, equivalently $\chi_{t}=0$). Thus, repulsions are exclusively short-ranged via Eq.(2) which prevents overlap of particles. First, we vary $\Lambda_{r}$ and $\phi$ and determine the polar order of the system. The phase diagram delimiting the CLF phases (polar flocks, bands, and disordered states) are shown in Figure 2(a). The states are delimited via the mean global polarization of the system at the steady state, $P_{\mathrm{ss}}=\langle|\mathbf{P}|\rangle_{\mathrm{ss}}$, where the polarization $\mathbf{P}$ is defined as the average of sum of the orientation of all the particles:
@@
The different phases shown in the phase diagram are distinguished with different values of the magnitude of $P_{\mathrm{ss}}$: flock ($P_{\mathrm{ss}}>0.9$), bands ($0.1<P_{\mathrm{ss}}<0.9$) and disordered ($P_{\mathrm{ss}}<0.1$). The snapshots of these three phases are presented in Fig. 2(b)-(d) respectively. The dynamics of $|\mathbf{P}|$ of the respective cases is shown in 2(e). On the $(\phi,\Lambda_{r})$ plane (Fig. 2(a)), we see that low-density flocks are sustained, that are eventually destabilized at sufficiently high repulsion and densities. These flocks have evidently a liquid-like spatial structure (Fig. 2(b)) [25, 43, 44]. The density field $\rho$ is thus not homogeneous across the system (see Fig. 2(b)). Density bands often occur in these systems, where particles move perpendicular to the length of these bands and travel Fig. 2(c)). For larger systems, many of these dense bands are observed over the space. See Movie I [45] for dynamics to obtain CLF starting with a disordered state.
III.1 Mechanism for the formation of CLF
The CLF flock forming region on Fig. 2(a) can be understood by consideration of competing effects between various length (time) scales of the problem. To this end, it is useful to consider the idealization of pair colloidal collision, where the pair collides move along a co-moving frame a distance $l_{sl}$ within a time $\tau_{sl}$, whilst rotating away (chemorepulsive). This can be interpreted as a sliding length and time scales respectively, and can be written as:
$\tau_{sl}=\frac{b}{v_{0}\Lambda_{r}},\qquad l_{sl}=v_{0}\tau_{sl}=\frac{b}{\Lambda_{r}}.$ (7)
Let us denote $\Lambda_{r}^{-}$ as the smallest value of $\Lambda_{r}$ to sustain the CLF (corresponding to the lowest transition line in Fig.2a), and $\Lambda_{r}^{+}$ as the largest value of $\Lambda_{r}$ above which the CLF is not sustained. $\Lambda_{r}^{+}$ is density dependent, as will be described below. The mechanism for the CLF can then be understood as a competition between $\Lambda_{r}$ and $\Lambda_{r}^{+}$, $\Lambda_{r}^{-}$; this is illustrated in Fig. 2 (f-h) (annotated accordingly in the phase diagram of Fig.2(a)). From (7), the mechanism can be understood in terms of sliding time scale $\tau_{sl}$ along with $\tau_{sl}^{-}$ and $\tau_{sl}^{+}$. There are three distinct cases (see corresponding label on phase diagram in Fig. 2(a)):
Case I. When $\Lambda_{r}<\Lambda_{r}^{-}$ (corresponding to $\tau_{sl}>\tau_{sl}^{-}$), the orientational changes of the colloids are too slow compared to the free movement time, pair colloids slide excessively long, hence local polar order cannot be sustained (Fig. 2 (f)).
Case II. In the regime of $\Lambda_{r}^{-}<\Lambda_{r}<\Lambda_{r}^{+}$ (equivalently $\tau_{sl}^{+}<\tau_{sl}<\tau_{sl}^{-}$), we have that the chemical response time fast enough such that local polar order is sustained, but still sufficiently slow such that the forward-backward symmetry at the pair collision level is broken (Fig. 2 (g)).
Case III. Finally, for $\Lambda_{r}>\Lambda_{r}^{+}$ (equivalently $\tau_{sl}<\tau^{+}_{sl}$), the chemical response is sufficiently quick such that there is no symmetry breaking at the pair collision level, and hence no local alignment (See Fig. 2 (h)).
III.2 Flocking by sliding an optimal distance while turning away
Using Eq.(7), we may also describe the mechanism for CLF in terms of sliding length scale $l_{sl}$, corresponding to $\tau^{+}_{sl}$ and $\tau^{-}_{sl}$. These are:
$l_{sl}^{+}=v_{0}\tau^{+}_{sl},\qquad l_{sl}^{-}=v_{0}\tau^{-}_{sl}.$
Note that in this choice of notations, $l_{sl}^{-}>l_{sl}^{+}$ as the time scale $\tau^{-}_{sl}>\tau^{+}_{sl}$. From simulations at $\phi\approx 0.31$, we find these values to be $l_{sl}^{+}\sim 8.3b$ ($\Lambda_{r}^{+}\sim 0.12$) and $l_{sl}^{-}\sim 100b$ ($\Lambda_{r}^{-}\sim 0.01$) , thus for low densities $l_{sl}^{+}$ of a few times the colloidal radius is feasible, whereas the upper limit spans the box size, which prohibits any flock formation. As we will show (and derive) below, $l_{sl}^{+}$ is directly proportional to density (with $l_{sl}^{-}$ fixed by the system size), thus CLF formation can be equivalently understood by the competing effects of density-dependent inter-particle (pair) collisions and sliding. We see that the phase boundary in Fig. 2(a) takes the functional form $\Lambda_{r}^{+}={\alpha}/{\rho}$. Here, $\alpha$ is constant that we will later quantify and derive. For densities above a critical density $\rho_{c}=(\frac{\alpha}{b})^{2/3}\approx 0.05$, $l_{sl}^{+}$ always exceeds the mean free path $\frac{1}{\sqrt{\rho}}$ so that flock formation can happen uninhibited. Indeed for very dilute systems ($\rho<0.05$, smaller than the densities displayed in Fig. 2), there is no flock formation as there is a large free path movement and thus pair sliding is infrequent (not shown here). For high densities, though pair collisions are more frequent, $l_{sl}>l_{sl}^{+}$ cannot be satisfied. Indeed, at these densities, $l_{sl}^{+}$ and $l_{sl}^{-}$ converge and flock formation is prohibited. Thus, for a finite range of densities the CLF can form and stably persist. Equivalently, there is no CLF if the pair colloids slide is too short ($l_{sl}<l_{sl}^{+}$) or too long ($l_{sl}>l_{sl}^{-}$).
We clarify at this juncture that such a pair collision mechanism is constructed as an idealization; in actual simulation, pair colloids may not slide an entire $l_{sl}$ before leaving a given collision radius. However, such an abstraction helps us understand the polar alignment mechanism at the pair colloidal level, and in addition agrees fully with simulation results at the collective level. A key feature of this mechanism to be emphasized is the requirement of short-ranged repulsion (provided here via excluded volume interactions), without which CLF formation at $\chi_{t}=0$ would not occur; this instead would have to be supplemented with $\chi_{t}>0$ [26]. Thus, (i) short-ranged translational repulsion and (ii) long-range orientational torques are the minimal requisite ingredients required for the formation of CLF.
Figure 3: (a) Phase diagram in the $(\Lambda_{r},\Lambda_{t})$ plane for the case of $\chi_{t}>0$. Dashed line refers to lower bound on $\chi_{r}$, solid line for lower bound on $\chi_{r}$ from Eq. (8). (b) shows a snapshot of the CCF phase (same color scheme to indicate the orientation of the particles as in Fig (2)).
IV CCF via chemo-repulsive forces and torques
@@
We now switch on $\chi_{t}$ and ask how the properties of the flock differ. On the $(\Lambda_{r},\Lambda_{t})$ plane, we find that there is a clear region of flocking for non-zero $\chi_{t}$ (see Figure 3(a)). We further find that the flocks have crystalline order (see Figure 3(b)). Such crystalline flocks in repulsive systems have been recently reported [26]. Similarly to the CLF, the CCF destabilized by sufficiently high rotational torques. In addition, $\Lambda_{r}^{+}$ now is linear in $\Lambda_{t}$. $\Lambda_{r}^{+}$ being larger for a positive $\Lambda_{t}$, this thus renders a shorter sliding length for pair colloids during the CCF formation. The net result is thus a shorter sliding length compared to the CLF, which we can attribute to the difference between crystalline and liquid structures. See Movie II [45] for dynamics to obtain CCF starting with a disordered state. We note that if this repulsion is too strong, the CCF is destabilized to a disordered gas (not shown in Figure 3).
The phase diagrams for the collective dynamics of the system have been presented in Figures (2)(a) and (3)(a) in terms of the three key dimensionless parameters: $\Lambda_{t}$, $\Lambda_{r}$ and $\phi$. The most generic form of the transition lines from the phase diagrams can be written as:
$\Lambda_{r}<\Lambda_{r}^{+},\qquad\Lambda_{r}^{+}=\frac{\alpha}{\rho}+{\beta\Lambda_{t}}.$ (8)
We obtain $\beta$ and $\alpha$ by first fitting the boundary line in Fig. 3(a). The slope gives $\beta\approx 0.167$, whilst the $x$-intercept gives $\alpha\approx 0.013/b^{3}$. This fit is obtained by first applying determining points of greatest variance of the order parameter (susceptibility) on the phase plane, and then applying least-squares regression on those points. The fitted $\alpha$ thus independently fixes the boundary on Fig. 2(a). The term $\alpha/\rho$ inhibits sliding at high densities. The second term due to the presence of long-ranged translational repulsion suppresses the sliding length and contributes to the crystalline structure. The delimiting line of $\Lambda_{r}^{-}$ on the left (vertical) of Fig. 3(a) and bottom (horizontal) of Fig. 2(a) are equal. Thus, the CCF mechanism may be summarized by being an extension of the CLF mechanism, with the additional long-ranged repulsive forces between the particles rendering the crystalline structure.
V Differences between CCF and CLF
Figure 4: (a) and (b) compare $g(r,\varphi)$ for the case of the CLF and the CCF. In (c), a comparison of the flocking phase, for CLF (red) and CCF (blue), is shown for the pair correlation $g(r)$.
V.1 Pair correlation function
The difference in spatial structure between the CLF and CCF is readily distinguished by the pair correlation function, which is given by:
$g(r)=\frac{1}{N}\sum_{i,j}\langle\delta\left(|\mathbf{r}_{i}-\mathbf{r}_{j}|-r\right)\rangle$ (9)
@@ V.1 Pair correlation function
This is plotted in Fig. (4)(c), distinguishing the solid and liquid signatures [25]. A closely related quantity is the polar pair correlation $g(r,\varphi)$ which explicitly depends on the orientations of the colloids. Here, $\cos(\varphi_{ij})=\mathbf{e}_{i}\cdot\hat{\mathbf{r}}_{ij}$, where $\hat{\mathbf{r}}_{ij}=(\mathbf{r}_{j}-{\mathbf{r}}_{j})/|\mathbf{r}_{j}-{\mathbf{r}}_{j}|$[46]. For the CLF, we see that there are quadrants where neighbours are more likely to be found (4(a)), as opposed to the hexatic order of the CCF (4(b)). In the case of CCF, the intensity (proportional to the probability of finding the particle from the reference - see Appendix A) in the well-localized hexagonal positions of the crystal are substantially higher than in the CLF, which are instead less intense with a greater relative spread. The four-lobe distribution is indicative of a highly anisotropic liquid phase where there is a decreased probability of finding a neighboring particle either at the front or back of the reference particle due to mutual repulsion.
@@
Figure 5: Phase diagrams in terms of (left) global hexatic order ($\psi_{6}$) and (right) density variance. The top panel (a) show these for the CLF, whereas the bottom panel (b) for the CCF. Dashed curves in each plot indicate order-disorder (polarization) transition line as in Figures 2(a) and 3(a).
V.2 Hexatic order and density variance
@@
Here, $\theta_{ij}$ is the angle between particle $i$ and particle $j$, and $N^{n}_{i}$} is the number of neighbors for particle $i$ ($N^{n}\approx 6$ for all data points). The observable $\psi_{6}$ is then averaged over $1000$ different particle configurations at the steady-state. We plot $\psi_{6}$ for the two cases ($\chi_{t}=0$ and $\chi_{t}>0$) in the left panel of Fig. 5. The transition lines from the phase diagrams in Figures 2(a) and 3(a) are also shown as dotted lines for guide to the eye. In the case of $\chi_{t}=0$ (Fig. 5(a) left), the hexatic order exists in the flocking phase, only for a very high density (area fraction $\phi=\rho\pi b^{2}>0.6$. Whereas, in case of $\chi_{t}>0$ (Fig. 5(b) left), with $\phi=\rho\pi b^{2}=0.36$, the system becomes highly hexatic in the flocking phase. However, near the transition line, where density bands are prone to occur, we find that the hexatic order gradually decreases.
@@
Next, density variance for the two cases ($\chi_{t}=0$ and $\chi_{t}>0$) are shown in the right panel of Fig. 5. The density variance is calculated from the density distribution for a specific parameter set shown in the phase diagram. For local density measurements, the Voronoi tessellation method has been implemented for a spatial configuration of particles. We calculate local density $\rho_{l}={\pi b^{2}}/{A_{l}}$, where $A_{l}$ is local area assigned to each particle. Then we compute the density variance as $\langle(\Delta\rho)^{2}\rangle=\langle\rho_{l}^{2}\rangle-\langle\rho_{l}\rangle^{2}$ (spatial average) and finally averaged over $1000$ configurations. In both cases, the density variance is much lower in the flocking state. However, it shows an increase near the transition region. The density variance is high for the high area fraction ($\phi=\rho\pi b^{2}$) and in the disordered phase, for the case $\chi_{t}=0$ (Fig. 5(a) right panel). For $\chi_{t}>0$ case, with low $\chi_{t}$ values, the variance is more prominent, shown in Fig. 5(b) (right panel).
Figure 6: Finite-size effects of phase transition. Top panel : (a) shows magnitude of the polar order parameter $|\mathbf{P}|$ versus $\Lambda_{r}$ for the CLF ($\Lambda_{t}=0$) for different area fractions $\phi$, whilst (b) displays these for different system of sizes $L$. Bottom panel: similarly plotted for the CCF, with the variation with respect to $\Lambda_{t}$ instead for different $\phi$ (c) and $L$ (d) respectively. Here, we have fixed $\Lambda_{r}=0.55$. The total number of particles NN corresponding to the $L$ are given as: ($L=50,N=239$), ($L=100,N=955$), ($L=200,N=3820$) and ($L=400,N=15279$).
V.3 Phase transition and finite size effects